Fermi National Laboratory
 

Accelerator Update for April 20, 2007 - April 23, 2007

 

Friday April 20

The day shift began with Operations monitoring the Recycler stash (227.2E10) and the Antiproton Source (Pbar) stack (71.6E10), with the Linac down due to problems with an RF station (LRF2), with the Booster, Main Injector (MI), NuMI, and MiniBooNE all down due to Linac problems, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

Linac technicians completed their work on LRF2. Operations sent beam to the Linac absorber at 9:20 AM.

Operations established beam to MiniBooNE at 9:27 AM.

At 9:36 AM, experts began work on the MI RF system. They completed their work at 10:01 AM.

At 10:20 AM, LRF2 tripped off. Experts investigated.

At 11 AM, Operations reported that Linac techs would make extensive repairs on LRF2. During this work, cryo system techs would replace the flywheel for a Tevatron (TeV) sector A3 wet engine.

Pbar experts began a 50-minute stacktail study at 11:20 AM.

LRF2 work was completed by 1:27 PM.

Operations established beam to MiniBooNE and NuMI at 1:36 PM.

At 1:43 PM, a Linac Klystron station (KRF3) tripped off. It reset.

Operations reported at 2:16 PM that Pbar experts had completed work on the lithium lens' power supply fans.

Operations established beam to Meson MTest experiment at 3:25 PM.

Operations reported at 4:20 PM that the Recycler had stashed antiprotons once during the day shift.

Operations reported at 5:14 PM that the TeV was ramping.

Operations put the TeV through a dry squeeze at 5:36 PM.

Operations put the TeV through a 4x0 wet squeeze at 7:03 PM.

A Cryo system expert reported at 7:18 that there appeared to be a bypass flow restriction in the TeV sector C4 leads. Experts made adjustments for the restriction.

At 7:22 PM, the I- Source vacuum gauge tripped off. It reset.

At 7:26 PM, there was a TeV sector B17 separator spark.

Operations put the TeV into shot setup at 7:35 PM.

Operations established store 5366 at 9:33 PM with an initial luminosity of 249.1E30.

At 11:03 PM, Operations reported MI beam position problems. An expert recommended adjustments.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Saturday April 21

The midnight shift began with Operations monitoring store 5366 (154.47E30) and the stash (93.6E10), with Pbar stacking (15.5E10), and with NuMI and MiniBooNE taking beam.

At 12:08 AM, KRF4 tripped off. It reset.

At 12:59 AM, Operations reported that a Booster RF station (BRF6) had tripped off. It reset.

At 1:09 AM, the TeV suffered at sector F4 quench possibly due to a B11 separator spark. Operators put the separators into a conditioning mode.

At 1:22 AM, Operations reported the first of several MI trips due to bad beam positions and power supply trips. Operations investigated and discovered that the trips were due to MECAR taking in some bad information. Operators cleared this.

At 4:20 AM, MI-20 suffered several power supply trips. This ended up being due to another MECAR problem.

Operations established beam to MTest at 6:07 AM.

Operations took the TeV separators out of conditioning mode at 7:01 AM.

Operations put the TeV through a 4x0 wet squeeze at 7:41 AM.

At 7:51 AM, a TeV RF station (TRF3) tripped off during the wet squeeze. Operations contacted an RF expert.

Operations reported at 8:37 AM that the Recycler had stashed twice during the midnight shift.

At 8:21 AM, LRF3 tripped off for the first of three times in about 90 minutes. It reset.

Operations established store 5368 at 11:08 AM with an initial luminosity of 177.1E30.

A MI expert began a study period at 2:32 PM.

At 5:32 PM, Operations reported that the Pelletron had tripped 12 times due to MI beam position problems.

At 5:47 PM, Operations reported a TeV sector D17 separator spark that caused a small loss of luminosity.

At 6:22 PM, the Pbar Debuncher tripped off due to a gradient regulation problem. It reset.

The MTest experiment reported that they were done with beam for the night at 6:35 PM.

Operations reported at 7:57 PM that the Recycler had stashed four times during the last 12 hours. Accelerator status: store 5368 was at 61.01E30, the stack was at 9.2E10, and the stash was at 164.5E10.

At 8:07 PM, KRF4 tripped off. It reset.

At 8:35 PM, Operations reported that MECAR had again learned in some bad data and had to be reset.

At 10:25 PM KRF7 tripped off on a LCW flow trip. It reset.

At 10:57 PM, the Tevatron Electron Lens (TEL1) tripped off due to a vacuum interlock. It reset.

Sunday April 22

The midnight shift began with Operations monitoring store 5368 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations established beam to MTest at 4:02 AM.

At 5:56 AM, KRF4 tripped off and with some operator help resumed normal running.

Operations reported at 8:01 AM that the Recycler had stashed five times during the evening shift.

Operations terminated store 5368 at 8:04 AM.

Operations put the TeV into shot setup around 8:18 AM.

At 10:06 AM, KRF4 tripped off. It reset.

Operations established store 5369 at 10:12 AM with an initial luminosity of 253.69E30.

At 10:41 AM, BRF11 tripped off. It reset.

At 11:35 AM, KRF3 tripped off. It reset.

At 12:11 PM, KRF4 tripped off. It reset.

At 3:37 PM, NuMI tripped off due to a bad beam position. A MI Lambertson (LAM61) also tripped off and wouldn’t reset. Operations called in an expert who had the Lambertson up and running by 4:38 PM. Operations resumed sending beam to Pbar, NuMI, and MiniBooNE.

At 5:57 PM, Operations reported that they were still having problems with the Pelletron. They contacted experts.

MTest signed off for the night at 7:05 PM.

Operations reported at 7:59 PM that the Recycler had stashed four times during the last 12 hours. Accelerator status: store 5369 was at 84.92E30, the stack was at 60.5E10, and the stash was at 122.9E10.

At 9:28 PM, Operations reported that the Pbar lithium lens had tripped off due to a power supply over temperature. Operators turned on a fan and tried to reset it, but the power supply would not reset. They contacted experts who suggested a Time Line Generator (TLG) event change. This resolved the problem.

At 9:55 PM, LRF3 tripped off for the first of two times in 9 minutes. It reset.

At 10:25 PM, KRF4 tripped off. It reset.

Monday April 23

The midnight shift began Operations monitoring store 5369 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 12:06 AM, LRF3 tripped off for the first of two times in 23 minutes. It reset.

At 12:58 AM, KRF3 tripped off. It reset.

At 3:20 AM, Operations reported more MECAR problems. Operations contacted an expert who suggested a CAMAC crate power cycle.

At 3:40 AM, the Pbar lithium lens tripped off. It reset.

At 3:45 AM, LRF3 tripped off for the first of three times in about 45 minutes. It reset.

MTest resumed taking beam at 5:16 AM.

At 5:23 AM, KRF4 tripped off. It reset.

At 7:09 AM, Operations reported the MECAR again learned in some bad information. Operators reset it.

Operations reported at 7:54 AM that the Recycler had stashed three times during the last 12 hours.

The Plans for Monday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/23/2007    email Fermilab